OpenStudy (anonymous):

How would I write the square root of 64z^4 in the calculator?

OpenStudy (rmrjr22):

\[\sqrt{64z^4}\]

OpenStudy (anonymous):

Yes!

OpenStudy (rmrjr22):

or \[(64z^4)^{1/2}\]

OpenStudy (anonymous):

You had it the first time, how do I calculate this?

OpenStudy (rmrjr22):

whats 8^2?

OpenStudy (anonymous):

64

OpenStudy (rmrjr22):

thats the same as \[\sqrt{64}\]

OpenStudy (rmrjr22):

\[\sqrt{z^4}= z^2\]

OpenStudy (anonymous):

What about z ^4

OpenStudy (rmrjr22):

= 8z^2

OpenStudy (anonymous):

I got that answer but thought that I was wrong. Thank you!

OpenStudy (rmrjr22):

welcome